# EvolveSignal: A Large Language Model Powered Coding Agent for Discovering Traffic Signal Control Algorithms

### v2

@@ Abstract
In traffic engineering, the fixed-time traffic signal control remains widely used for its low cost, stability, and interpretability. However, its design depends on hand-crafted formulas (e.g., Webster) and manual re-timing by engineers to adapt to demand changes, which is labor-intensive and often yields suboptimal results under heterogeneous or congested conditions. This paper introduces the EvolveSignal, a large language models (LLMs) powered coding agent to automatically discover new traffic signal control algorithms. We formulate the problem as program synthesis, where candidate algorithms are represented as Python functions with fixed input–output structures, and iteratively optimized through external evaluations (e.g., a traffic simulator) and evolutionary search. Experiments on a signalized intersection demonstrate that the discovered algorithms outperform Webster’s baseline, reducing average delay by 20.1% and average stops by 47.1%. Beyond performance, ablation and incremental analyses reveal that EvolveSignal modifications—such as adjusting cycle length bounds, incorporating right-turn demand, and rescaling green allocations—can offer practically meaningful insights for traffic engineers. This work opens a new research direction by leveraging AI for algorithm design in traffic signal control, bridging program synthesis with transportation engineering.
Index Terms:
@@
I Introduction
Traffic signal control remains the cost-effective strategy for mitigating urban congestion and improving traffic efficiency. In recent years, artificial intelligence (AI)-based methods have attracted significant attention, especially those utilizing reinforcement learning (RL) [1, 2] and large language models (LLMs) [3, 4]. These data-driven approaches demonstrate the potential to learn high-performance control strategies that can enhance mobility at both the intersection and network levels.
@@ I Introduction
In contrast, fixed-time traffic signal control is still widely adopted due to its simplicity, low implementation cost, and operational stability [5, 6]. More importantly, fixed-time control methods are inherently interpretable and transparent, making them easier to adjust and integrate with domain-specific rules. Optimizing fixed-time control plans remains a practically valuable and promising research direction, especially in scenarios where sensing and actuation resources are limited.
@@
Traditional fixed-time control algorithms are typically derived from empirical engineering heuristics, supported by traffic flow theory, queuing models, and optimization methods. In this study, we define a fixed-time traffic signal control algorithm as a method that computes a static signal timing plan based on low-resolution or aggregated traffic demand data, without requiring real-time sensing. The output typically includes a fixed cycle length and green time allocation. Among these, Webster’s method [7] is a classic example that uses critical flow ratios and cycle losses to derive closed-form solutions.
@@
While these classical methods are highly interpretable and easy to implement, they are often insufficient to meet the specific operational needs of individual intersections. Due to heterogeneity in traffic demand, geometry, and performance objectives across intersections, a single globally applied strategy (e.g., directly using Webster’s output) may lead to suboptimal outcomes. In practice, traffic engineers are frequently required to fine-tune or adapt these algorithms on a per-intersection basis—adjusting parameters, modifying constraints, or integrating additional domain rules—to better match local conditions and achieve desired performance.
@@
This leads to a critical research question: Can we move beyond hand-crafted rules and enable automated discovery of fixed-time control algorithms with higher adaptability and performance?
To address this, we propose leveraging the algorithmic discovery and code generation capabilities of LLMs to optimize fixed-time control algorithms in code (e.g., Python). Specifically, we introduce a novel framework, EvolveSignal, which treats a fixed-time signal control algorithm as a function-level program (e.g., a Python function with fixed inputs and outputs) and aims to optimize its internal computation logic—how input variables are processed, intermediate quantities are combined, and outputs are derived. The main contributions of this paper are as follows:
@@
  • We formulate fixed-time signal control as a program synthesis problem, representing an algorithm as a symbolic function with fixed input–output structure. The focus is on discovering new algorithmic logic, how traffic features are transformed into signal timings.
@@
  • We develop EvolveSignal, an LLM-powered coding agent that explores, generates, and evaluates candidate programs through simulation-based evaluation and evolutionary search. The framework iteratively refines executable code, enabling interpretable and auditable algorithm discovery.
@@
  • We validate the EvolveSignal on synthetic traffic scenarios with heavy traffic conditions, showing significant improvements over Webster’s baseline in terms of delay reduction and stop minimization, while revealing interpretable modifications that can provide useful insights for traffic engineering practice.
@@
The remainder of this paper is organized as follows. Section II reviews related work on conventional and AI-based traffic signal control. Section III introduces the proposed EvolveSignal framework. Section IV presents the experimental setup, results, and ablation analyses. Section VI discusses insights and implications. Section VII concludes the paper and outlines future directions.
II Related Work
II-A Conventional traffic signal control methods
@@
Traditional traffic signal control methods for a single intersection can be divided into three groups. The first is fixed-time signal control, such as Webster’s method, which optimizes green time using historical traffic data to minimize delay [7]. The second is actuated signal control, where phases are adjusted based on real-time data from detectors [8]. Both methods, however, struggle to handle dynamic traffic patterns. A third approach, adaptive traffic signal control, dynamically adjusts phases and green times based on real-time conditions [9].
@@
From a control scope perspective, traditional methods can also be classified into single-intersection control, arterial coordination [10, 11], and network-wide control [12]. While these methods have a solid theoretical foundation and are relatively easy to implement, they often require the traffic management engineer to optimize each intersection individually based on local conditions, which is time-consuming, labor-intensive, and costly.
II-B AI-based traffic signal control methods
@@
Recently, RL-based approaches have shown significant promise in adaptive traffic signal control, achieving superior performance in terms of reduced delay [13, 2], improved energy efficiency [14, 15], and enhanced safety [16] compared to traditional methods. In parallel, LLM-based methods have shown unique advantages across various transportation domains [17, 18, 19, 20]. In the context of adaptive traffic signal control, LLMs can deliver even better adaptive control performance while providing natural language reasoning and explanations[3, 4]. However, practical deployment remains challenging due to high costs, privacy concerns, and the complex infrastructure required to scale RL and LLM models. This study proposes a novel approach to better harness the potential of AI for traffic signal control, ensuring both scalability and real-world feasibility.
III Methodology
III-A Problem Definition
@@
where X{X} denotes the input space, including structural and traffic-related features of an intersection (e.g., traffic volume, geometric layout, and phase configuration); Y{Y} represents the output space of fixed-time signal control parameters (e.g., cycle length, green time per phase); and $f_{\theta}$ is a parameterized executable program that encodes the logic and structure of the control algorithm using symbolic elements such as variables, operators, formula templates, and conditional statements, as shown in Fig. 1.
@@
III-B Overview of the Proposed Framework
Fig. 2 shows the overall framework of EvolveSignal, an LLM-powered coding agent designed to evolve fixed-time traffic signal control algorithms. The process begins with an initial program (e.g., an implementation of Webster’s method), which is iteratively modified, simulated, and evaluated until convergence. In each iteration, candidate programs are stored and organized in the Program Database, which supports both performance tracking and diversity maintenance. A parent program and several inspiration programs are then passed to the Prompt Sampler, which constructs prompts combining code, metrics, and search strategies. These prompts are processed by the LLMs Ensemble, generating code-level modifications that act as crossover and mutation operators. The resulting child program is tested in the Evaluators Pool using microscopic traffic simulation, and its performance is fed back to the Program Database. This loop continues until either the maximum number of iterations is reached or the best-performing algorithm is identified. Detailed descriptions of these modules, as well as example prompts and LLM responses, are provided in supplementary Appendices A and B.
@@ III-B Overview of the Proposed Framework
III-C Pseudocode for the Proposed Framework
The workflow of EvolveSignal is summarized in Algorithm 1, adapted from the AlphaEvolve framework [21] with implementation modifications from the open-source OpenEvolve project [22]. The pseudocode captures the evolutionary loop of selection, LLM-guided modification, simulation-based evaluation, and population update.
@@ III-C Pseudocode for the Proposed Framework
11:  Evaluate $f_{\theta}^{\prime}$ in SUMO [23], compute performance score $\text{Score}(f_{\theta}^{\prime})$
12:  Update Program Database with $f_{\theta}^{\prime}$ using MAP-Elites [24]
@@
IV Experiment
IV-A Experimental Setup
Experiments were conducted on a simulated four-leg intersection in SUMO (Fig. 3). Each approach is 400 m long with four lanes: one left-turn, two through, and one shared through–right lane. The yellow interval is 3 s and the all-red interval is 1 s. Minimum greens are 20 s for through and 15 s for left turns.
@@ IV-A Experimental Setup
The goal is to evaluate whether the proposed method improves the algorithm’s performance under heavy demand. The intersection operates with four fixed phases: east-west (EW) through, EW left-turn, north-south (NS) through, and NS left-turn phases. Three scenarios (S1–S3) represent distinct demand patterns with overall heavy traffic conditions. Scenario S1 is balanced; S2 emphasizes NS flows; S3 emphasizes EW through and NS left. Demands and Critical Ratio Sum (CRS) values are listed in Table I.
TABLE I: Traffic demand scenarios and Critical Ratio Sum (CRS). All values are in veh/h, formatted as through/left/right.
@@
Webster’s Python implementation serves as the initial program. The LLMs Ensemble samples from DeepSeek-v3 [25] (40%), DeepSeek-r1 [26] (10%), OpenAI-o4-mini-high (40%), and OpenAI-o3 (10%), with temperature 0.6.
IV-B Evaluation Metrics
@@
V Experiment Results
Fig. 4 shows the evolution path of the best discovered program after 300 iterations, along with its performance. The performance generally shows an upward trend with occasional plateaus or even declines during some iterations. Overall, as shown in Table II, the combined score improved from 0.4893 (for the initial program) to 0.5946, which represents a 21.5% improvement. The optimized program outperforms the initial program with a 20.1% reduction in average delay and a 47.1% reduction in the number of stops, demonstrating the effectiveness of the evolutionary approach.
@@ V Experiment Results
TABLE II: Performance Comparison Between the Initial Program and the Best Discovered Program After 300 Iterations
@@
Fig. 8 and Fig. 9 present the complete Python code of the initial program and the best discovered program. The discovered program introduces several key modifications (highlighted by blue dashed lines in Fig. 9), with the underlying reasoning provided by the LLMs:
  • Cycle Length Bound (CLB). The maximum cycle length was increased from 130 s to 240 s. As shown in Fig. 10, the LLM argued that a longer cycle allows more effective green time in high-demand scenarios, thereby reducing delay and stops.
  • Right-Turn Inclusion (RTI). Right-turn flows were added to the through demand calculation. Fig. 11 illustrates how the LLM observed that right-turn vehicles were omitted in the initial program. Given the presence of through–right shared lanes, it is suggested to include right-turn flows to better balance capacity.
  • Shared Lane Factor (SLF). The contribution of the shared through–right lane was reduced from 0.9 to 0.5. As shown in Fig. 12, the LLM reasoned that a factor of 0.9 overestimated lane capacity, which underestimated the required green time for through movements. Reducing the factor to 0.5 raised their priority, allocating more green to critical through directions.
  • Minimum-Green Feasibility (MGF). The cycle length was adjusted when the effective green was insufficient to satisfy mandatory minimums. As noted by the LLM (see Fig. 13), the initial thresholds (50, 90) were too short to cover the required minimums (20 s per through phase and 15 s per left-turn, plus 16 s for intergreens, totaling 86 s). It accordingly revised the range to (90, 140) to ensure feasibility.
  • Post-Allocation Rescaling (PAR). Green times were proportionally rescaled to match the effective green budget. As noted by the LLM (see Fig. 11) this avoided both unused “dead” green seconds, which increase delay, and overruns, which reduce capacity for other phases.
@@
Table III shows the ablation results when removing individual modifications from the discovered program. Among all components, CLB is the dominant contributor: its removal decreases $S_{c}$ from $0.5946$ to $0.4812$ ($19.07\%$). The benefit of CLB stems from longer cycles diluting the fixed 16 s lost time, thereby increasing the share of effective green under heavy demand. Crucially, the improvement does not arise from arbitrarily lengthening cycles but from avoiding performance degradation due to overly restrictive bounds. RTI and SLF provide secondary yet important gains: removing either reduces $S_{c}$ by $3.48\%$ and $2.98\%$, respectively, reflecting their role in balancing green allocation by accounting for right-turn demand and correcting shared-lane capacity. PAR offers a modest but stable gain ($0.81\%$ when removed), largely due to rounding effects in high-demand scenarios; it would be more impactful under heterogeneous traffic when some phases approach minimum-green thresholds. Finally, MGF has no effect here since all tested cycles are sufficiently long, though it is expected to safeguard feasibility under lighter traffic.
@@
To further analyze the most effective strategies, Table IV presents an incremental analysis starting from the initial program. Again, CLB dominates: adding CLB alone improves $S_{c}$ by $15.63\%$. Adding SLF yields a minor gain ($1.51\%$), while adding RTI alone surprisingly degrades performance ($2.27\%$), as incorporating extra demand without cycle or capacity adjustment distorts green allocation. Notably, combining modifications yields synergistic rather than additive effects. For instance, CLB+RTI and CLB+SLF improve $S_{c}$ by $17.90\%$ and $17.29\%$, respectively, exceeding the sum of individual contributions. This suggests a coupling between demand modeling and cycle flexibility: RTI or SLF only becomes effective when longer cycles allow sufficient redistribution of green. These findings highlight two insights: (i) some modifications are detrimental in isolation, and (ii) evolutionary search is crucial for identifying synergistic combinations that outperform individual strategies. Beyond performance, this also illustrates how a coding-agent framework can uncover non-intuitive design patterns that may be difficult for human engineers to identify through manual, additive reasoning.
TABLE III: Ablation study of the optimized program by removing individual modifications. (“w/o” denotes removal of a component).
@@
TABLE IV: Incremental analysis by progressively adding modifications to the initial program. (“+” denotes addition of a component)
@@
VI Discussion
VI-A Types of LLM-Generated Modifications

### v3

# EvolveSignal: A Large Language Model Powered Coding Agent for Discovering Traffic Signal Control Strategies

@@ Abstract
In traffic engineering, fixed-time traffic signal control remains widely used for its low cost, stability, and interpretability. However, its design relies on hand-crafted formulas (e.g., Webster) and manual re-timing by engineers to adapt to demand changes, which is labor-intensive and often yields suboptimal results under heterogeneous or congested conditions. This paper introduces EvolveSignal, an LLM-powered coding agent for automatically discovering interpretable heuristic strategies for fixed-time traffic signal control. Rather than deriving entirely new analytical formulations, the proposed framework focuses on exploring code-level variations of existing control logic and identifying effective combinations of heuristic modifications. We formulate the problem as program synthesis, where candidate strategies are represented as Python functions with fixed input–output structures and iteratively optimized through external evaluations (e.g., a traffic simulator) and evolutionary search. Experiments on a signalized intersection demonstrate that the discovered strategies outperform a classical baseline (Webster’s method), reducing average delay by 20.1% and average stops by 47.1%. Beyond performance, ablation and incremental analyses reveal that EvolveSignal can identify meaningful modifications, such as adjusting cycle length bounds, incorporating right-turn demand, and rescaling green allocations, that provide useful insights for traffic engineers. This work highlights the potential of LLM-driven program synthesis for supporting interpretable and automated heuristic design in traffic signal control. The code for this paper is open-source and available at: https://github.com/georgewanglz2019/EvolveSignal.
keywords:
@@
journal: Elsevier
[…]
[a]organization=Department of Data Science and Artificial Intelligence, Monash University, city=Melbourne, country=Australia
[…]
[b]organization=School of Transportation, Southeast University, city=Nanjing, country=People’s Republic of China
[…]
[c]organization=R&D Center, Zhejiang Dahua Technology CO., LTD, city=Hangzhou, country=People’s Republic of China
[…]
[d]organization=Institute of Automation, Chinese Academy of Sciences, city=Beijing, country=People’s Republic of China
[…]
[e]organization=Department of Civil and Environmental Engineering, Monash University, city=Melbourne, country=Australia
[…]
[f]organization=Department of Civil and Architectural Engineering, KTH Royal Institute of Technology, city=Stockholm, country=Sweden
1 Introduction
Traffic signal control remains the cost-effective strategy for mitigating urban congestion and improving traffic efficiency. In recent years, artificial intelligence (AI)-based methods have attracted significant attention, especially those utilizing reinforcement learning (RL) (Wei et al., 2019; Wang et al., 2023) and large language models (LLMs) (Movahedi and Choi, 2024; Lai et al., 2025). These data-driven approaches demonstrate the potential to learn high-performance control strategies that can enhance mobility at both the intersection and network levels.
@@ 1 Introduction
In contrast, fixed-time traffic signal control is still widely adopted due to its simplicity, low implementation cost, and operational stability (Zhou et al., 2024; Wang et al., 2025b). More importantly, fixed-time control methods are inherently interpretable and transparent, making them easier to adjust and integrate with domain-specific rules. Optimizing fixed-time control plans remains a practically valuable and promising research direction, especially in scenarios where sensing and actuation resources are limited.
@@
Traditional fixed-time control algorithms are typically derived from empirical engineering heuristics, supported by traffic flow theory, queuing models, and optimization methods. In this study, we define a fixed-time traffic signal control algorithm as a method that computes a static signal timing plan based on low-resolution or aggregated traffic demand data, without requiring real-time sensing. The output typically includes a fixed cycle length and green time allocation. Among these, Webster’s method (Webster, 1958) is a classic example that uses critical flow ratios and cycle losses to derive closed-form solutions.
@@
While these classical methods are highly interpretable and easy to implement, they are often insufficient to meet the specific operational needs of individual intersections. Due to heterogeneity in traffic demand, geometry, and performance objectives across intersections, a single globally applied strategy (e.g., directly using Webster’s output) may lead to suboptimal outcomes. In practice, traffic engineers are frequently required to fine-tune or adapt these algorithms on a per-intersection basis by adjusting parameters, modifying constraints, or integrating additional domain rules to better match local conditions and achieve desired performance.
@@
This leads to a critical research question: Can we move beyond hand-crafted rules and enable automated discovery of effective heuristic strategies for fixed-time control with higher adaptability and performance?
To address this, we propose leveraging the code generation and reasoning capabilities of LLMs to explore variations of fixed-time control algorithms in executable form (e.g., Python functions). Rather than deriving entirely new analytical models, the goal is to automatically identify useful modifications and combinations of domain-inspired rules within an interpretable programming framework.
The main contributions of this paper are as follows:
  • 1.
@@
    We formulate fixed-time signal control as a program synthesis problem, where an algorithm is represented as an executable function with a fixed input–output structure. The focus is on discovering interpretable heuristic structures that transform traffic features into signal timings.
  • 2.
@@
    We develop EvolveSignal, an LLM-powered coding agent that explores and refines candidate programs through simulation-based evaluation and evolutionary search, enabling automated and auditable heuristic discovery.
  • 3.
@@
    We validate EvolveSignal on synthetic traffic scenarios with heavy demand, showing improvements over a classical baseline while revealing interpretable modifications that provide useful insights for traffic engineering practice.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work on conventional and AI-based traffic signal control. Section 3 introduces the proposed EvolveSignal framework. Section 4 presents the experimental setup, results, and ablation analyses. Section 5 discusses insights and implications. Section 6 concludes the paper and outlines future directions.
2 Related Work
2.1 Conventional traffic signal control methods
@@
Traditional traffic signal control methods for a single intersection can be divided into three groups. The first is fixed-time signal control, such as Webster’s method, which optimizes green time using historical traffic data to minimize delay (Webster, 1958). The second is actuated signal control, where phases are adjusted based on real-time data from detectors (Fellendorf, 1994). Both methods, however, struggle to handle dynamic traffic patterns. A third approach, adaptive traffic signal control, dynamically adjusts phases and green times based on real-time conditions (Wang et al., 2018; Xie et al., 2025).
@@
From a control scope perspective, traditional methods can also be classified into single-intersection control, arterial coordination (Gartner et al., 1991; Wang and Peng, 2022), and network-wide control (Hunt et al., 1982). While these methods have a solid theoretical foundation and are relatively easy to implement, they often require the traffic management engineer to optimize each intersection individually based on local conditions, which is time-consuming, labor-intensive, and costly.
2.2 AI-based traffic signal control methods
@@
Recently, RL-based approaches have shown significant promise in adaptive traffic signal control, achieving superior performance in terms of reduced delay (Chu et al., 2019; Wang et al., 2023; Kuwahara et al., 2025; Jin et al., 2026; Ren et al., 2025), improved energy efficiency (Koch et al., 2023; Peng et al., 2025), and enhanced safety (Gong et al., 2020) compared to traditional methods. In parallel, LLM-based methods have shown unique advantages across various transportation domains (Ma et al., 2026; Nie et al., 2025; Guo et al., 2024; Ling et al., 2026; Liu et al., 2025; Wang et al., 2025a; Pang et al., 2026; Qin et al., 2025a; Zhen and Yang, 2025; Qin et al., 2025b). In the context of adaptive traffic signal control, LLMs can deliver even better adaptive control performance while providing natural language reasoning and explanations(Movahedi and Choi, 2024; Lai et al., 2025). However, practical deployment remains challenging due to high costs, privacy concerns, and the complex infrastructure required to scale RL and LLM models. This study proposes a novel approach to better harness the potential of AI for traffic signal control, ensuring both scalability and real-world feasibility.
3 Methodology
3.1 Problem Definition
@@
where X{X} denotes the input space, including structural and traffic-related features of an intersection (e.g., traffic volume, geometric layout, and phase configuration); Y{Y} represents the output space of fixed-time signal control parameters (e.g., cycle length, green time per phase); and $f_{\theta}$ is a parameterized executable program that encodes the logic and structure of the control algorithm using symbolic elements such as variables, operators, formula templates, and conditional statements, as shown in Figure 1.
@@
3.2 Overview of the Proposed Framework
Figure 2 shows the overall framework of EvolveSignal, an LLM-powered coding agent designed to evolve fixed-time traffic signal control algorithms. The process begins with an initial program (e.g., an implementation of Webster’s method), which is iteratively modified, simulated, and evaluated until convergence. In each iteration, candidate programs are stored and organized in the Program Database, which supports both performance tracking and diversity maintenance. A parent program and several inspiration programs are then passed to the Prompt Sampler, which constructs prompts combining code, metrics, and search strategies. These prompts are processed by the LLMs Ensemble, generating code-level modifications that act as crossover and mutation operators. The resulting child program is tested in the Evaluators Pool using microscopic traffic simulation, and its performance is fed back to the Program Database. This loop continues until either the maximum number of iterations is reached or the best-performing algorithm is identified. Detailed descriptions of these modules are provided in Appendix 6.1.
@@ 3.2 Overview of the Proposed Framework
3.3 Pseudocode for the Proposed Framework
The workflow of EvolveSignal is summarized in Algorithm 1, adapted from the AlphaEvolve framework (Novikov et al., 2025) with implementation modifications from the open-source OpenEvolve project (Sharma, 2025). The pseudocode captures the evolutionary loop of selection, LLM-guided modification, simulation-based evaluation, and population update.
@@ 3.3 Pseudocode for the Proposed Framework
11:  Evaluate $f_{\theta}^{\prime}$ in SUMO (Lopez et al., 2018), compute performance score $\text{Score}(f_{\theta}^{\prime})$
12:  Update Program Database with $f_{\theta}^{\prime}$ using MAP-Elites (Mouret and Clune, 2015)
@@
4 Experiment
4.1 Experimental Setup
Experiments were conducted on a simulated four-leg intersection in SUMO (Figure 3). Each approach is 400 m long with four lanes: one left-turn, two through, and one shared through–right lane. The yellow interval is 3 s and the all-red interval is 1 s. Minimum greens are 20 s for through and 15 s for left turns.
@@ 4.1 Experimental Setup
The goal is to evaluate whether the proposed method improves the algorithm’s performance under heavy demand. The intersection operates with four fixed phases: east-west (EW) through, EW left-turn, north-south (NS) through, and NS left-turn phases. Three scenarios (S1–S3) represent distinct demand patterns with overall heavy traffic conditions. Scenario S1 is balanced; S2 emphasizes NS flows; S3 emphasizes EW through and NS left. Demands and Critical Ratio Sum (CRS) values are listed in Table 1.
Table 1: Traffic demand scenarios and Critical Ratio Sum (CRS). All values are in veh/h, formatted as through/left/right.
@@
Webster’s Python implementation serves as the initial program. The LLMs Ensemble samples from DeepSeek-v3 (Liu et al., 2024) (40%), DeepSeek-r1 (Guo et al., 2025) (10%), OpenAI-o4-mini-high (40%), and OpenAI-o3 (10%), with temperature 0.6.
4.2 Evaluation Metrics
@@
4.3 Experiment Results
It is important to note that the objective of this study is not to benchmark against all existing fixed-time optimization methods, but to evaluate whether LLM-guided program evolution can improve upon a classical and interpretable baseline. The use of Webster’s method as the initial program provides a transparent and widely recognized reference point for assessing improvements.
Figure 4 shows the evolution path of the best discovered program after 300 iterations, along with its performance. The performance generally shows an upward trend with occasional plateaus or even declines during some iterations. The initial performance drop reflects exploration in early iterations, where diverse but suboptimal candidate programs are generated before convergence toward higher-performing solutions. Overall, as shown in Table 2, the combined score improved from 0.4893 (for the initial program) to 0.5946, which represents a 21.5% improvement. The optimized program outperforms the initial program with a 20.1% reduction in average delay and a 47.1% reduction in the number of stops, demonstrating the effectiveness of the evolutionary approach.
@@ 4.3 Experiment Results
Table 2: Performance Comparison Between the Initial Program and the Best Discovered Program After 300 Iterations
@@
Figure 8 and Figure 9 present the complete Python code of the initial program and the best discovered program. The discovered program introduces several key modifications (highlighted by blue dashed lines in Figure 9), with the underlying reasoning provided by the LLMs:
  • 1.
    Cycle Length Bound (CLB). The maximum cycle length was increased from 130 s to 240 s. As shown in Figure 10, the LLM argued that a longer cycle allows more effective green time in high-demand scenarios, thereby reducing delay and stops.
  • 2.
    Right-Turn Inclusion (RTI). Right-turn flows were added to the through demand calculation. Figure 11 illustrates how the LLM observed that right-turn vehicles were omitted in the initial program. Given the presence of through–right shared lanes, it is suggested to include right-turn flows to better balance capacity.
  • 3.
    Shared Lane Factor (SLF). The contribution of the shared through–right lane was reduced from 0.9 to 0.5. As shown in Figure 12, the LLM reasoned that a factor of 0.9 overestimated lane capacity, which underestimated the required green time for through movements. Reducing the factor to 0.5 raised their priority, allocating more green to critical through directions.
  • 4.
    Minimum-Green Feasibility (MGF). The cycle length was adjusted when the effective green was insufficient to satisfy mandatory minimums. The LLM noted that the initial thresholds (50, 90) were too short to cover the required minimums (20 s per through phase and 15 s per left-turn, plus 16 s for intergreens, totaling 86 s). It accordingly revised the range to (90, 140) to ensure feasibility.
  • 5.
    Post-Allocation Rescaling (PAR). Green times were proportionally rescaled to match the effective green budget. As shown in Figure 11, this avoided both unused “dead” green seconds, which increase delay, and overruns, which reduce capacity for other phases.
@@
Table 3 shows the ablation results when removing individual modifications from the discovered program. Among all components, CLB is the dominant contributor: its removal decreases $S_{c}$ from $0.5946$ to $0.4812$ ($19.07\%$). The benefit of CLB stems from longer cycles diluting the fixed 16 s lost time, thereby increasing the share of effective green under heavy demand. Crucially, the improvement does not arise from arbitrarily lengthening cycles but from avoiding performance degradation due to overly restrictive bounds. RTI and SLF provide secondary yet important gains: removing either reduces $S_{c}$ by $3.48\%$ and $2.98\%$, respectively, reflecting their role in balancing green allocation by accounting for right-turn demand and correcting shared-lane capacity. PAR offers a modest but stable gain ($0.81\%$ when removed), largely due to rounding effects in high-demand scenarios; it would be more impactful under heterogeneous traffic when some phases approach minimum-green thresholds. Finally, MGF has no effect here since all tested cycles are sufficiently long, though it is expected to safeguard feasibility under lighter traffic.
@@
To further analyze the most effective strategies, Table 4 presents an incremental analysis starting from the initial program. Again, CLB dominates: adding CLB alone improves $S_{c}$ by $15.63\%$. Adding SLF yields a minor gain ($1.51\%$), while adding RTI alone surprisingly degrades performance ($2.27\%$), as incorporating extra demand without cycle or capacity adjustment distorts green allocation. Notably, combining modifications yields synergistic rather than additive effects. For instance, CLB+RTI and CLB+SLF improve $S_{c}$ by $17.90\%$ and $17.29\%$, respectively, exceeding the sum of individual contributions. This suggests a coupling between demand modeling and cycle flexibility: RTI or SLF only becomes effective when longer cycles allow sufficient redistribution of green. These findings highlight two insights: (i) some modifications are detrimental in isolation, and (ii) evolutionary search is crucial for identifying synergistic combinations that outperform individual strategies. Beyond performance, this also illustrates how a coding-agent framework can uncover non-intuitive design patterns that may be difficult for human engineers to identify through manual, additive reasoning.
Table 3: Ablation study of the optimized program by removing individual modifications. (“w/o” denotes removal of a component).
@@
Table 4: Incremental analysis by progressively adding modifications to the initial program. (“+” denotes addition of a component)
@@
5 Discussion
5.1 Types of LLM-Generated Modifications